A particle P of mass 5kg is suspended from a fixed point O by a light inextensible string of length 1m. The particle is projected from its lowest positon at the point A, with a horizontal speed of 4 ms-1. When angle AOP = 600. Find: a)The speed of P b)The tension in the string.


1
Expert's answer
2021-03-26T11:26:21-0400

Given:

mass (m)=5kg

length(l)=1m

angle(θ)=600

velocity(v)=4m\s


solution:


so according to problem we have to find tention(T) and speed (s) at 60o or point B .



a)from the figure we can calculate speed s by,

energy conservation


"{1\over2}mv^2={1\over2}ms^2+mg(l)(1-cos\theta)"


"{1\over2}(4)(5)^2={1\over2}(5)s^2+(5)(10)(1)({1\over2})"


"\boxed{s=\sqrt10m/s}"


b)using this we can calculate Tension by formula

"\boxed{T={mv^2\over l}+mgcos\theta}"


"T={5(10)\over 1}+(5)(10)({1\over2})"


"\boxed{T=75N}"